# Optimized Fish Locomotion using Design-by-Morphing and Bayesian Optimization

### v1

@@ 4.1 Performance Analysis of Optimal Profiles
To identify the optimal swimming profile, we employ the MixMOBO algorithm within a six-dimensional design space. The objective of the optimization is to maximize the propulsive efficiency $\eta$; equality reducing the work consumption by the foil to undulate.
@@
The optimization process begins with 50 randomly generated initial samples, followed by 150 optimization iterations (epochs) guided by the MixMOBO strategy. Figure 8 presents the evolution of the best-obtained efficiency over the 150 search iterations, along with the corresponding optimal profile shapes and undulation kinematics of the NACA 0012 airfoil. Details of some of the best-performing efficiency cases are also provided in Table 4, which reveals a remarkable efficiency range between 40–82%, which significantly exceeds that of the reference swimming modes, carangiform and anguilliform, within the domain of given kinematics parameters. Furthermore, the best-performing profiles demonstrate high efficiency at the lowest wavelength and highest frequency within the given range, while the conventional carangiform and anguiliform profiles attain their optimal performance predominantly at lower frequencies (see Fig. 7). To investigate the differences in the kinematic behavior between the optimized and conventional swimming modes, the backbone (centerline) undulation of the NACA 0012 airfoil over five different time instances within one oscillation cycle is presented in Fig. 9. The first and second optimal profiles exhibit similar undulatory patterns, with the main distinction being a more pronounced head motion in the optimal profile. Compared to the anguilliform profile, the optimal profile aligns closely over a significant region that spans approximately 30% to 87% of the body length. However, at the leading and trailing edges, notable differences are evident. The optimal profile exhibits head motion in the opposite direction to that of the anguilliform profile, and the tail amplitude is marginally lower, with the anguilliform tail reaching a maximum lateral amplitude near 0.1. These distinctive features, specifically the out-of-phase head motion, may contribute to the superior propulsive performance exhibited by the optimized profiles. The observed increase in efficiency is likely associated with these deviations from the anguilliform pattern in both head and tail kinematics. To further examine this hypothesis, a detailed analysis of the wake structures and force distributions is presented in the following discussion.
@@
Figure 8: Convergence history of MixMOBO showing the evolution of the best-obtained efficiency over 150 search iterations, together with the corresponding optimal profile shapes and undulation kinematics of the NACA 0012 airfoil.
@@
Design Epoch $\omega_{1}$ $\omega_{2}$ $\omega_{3}$ $\omega_{4}$ $\omega_{5}$ $\lambda^{*}$ $f$ $\eta$
3rd Best 13 -0.866 -0.250 0.108 0.0 0.375 0.5 4.5 0.428
2nd Best 17 -0.866 0.0 0.0 0.0 0.5 0.5 4.5 0.518
- - -0.866 0.0 0.125 0.0 0.433 0.5 4.0 0.651
Optimal 20 -0.866 0.0 0.125 0.0 0.433 0.5 4.5 0.824
@@
Figure 9: Time evolution of the backbone (centerline) kinematics of the top three optimized swimming profiles compared with the conventional anguilliform and carangiform modes at five different phases within one undulation cycle ($t/T=0.0$, $0.2$, $0.4$, $0.6$, and $0.8$). The optimized profiles exhibit notable differences in head and tail motions compared to the anguilliform profile.
Furthermore, in the optimal profile, a significant increase in efficiency from 65.1% to 82.4% is observed as the undulating frequency increases from 4.0 to 4.5. To quantify the influence of frequency on efficiency, flow simulations are performed for the top-performing profiles across a range of frequencies at fixed wavelengths of $\lambda^{*}=0.50$, $0.60$, and $0.70$. Figure 10 illustrates the mean forward (propulsive) velocity, work consumption, and propulsive efficiency of the top three performing profiles in comparison with the conventional anguilliform and carangiform swimming profiles, evaluated as functions of frequency for fixed wavelengths. The results indicate that the mean forward velocity consistently increases with frequency for all considered wavelengths. Although trends for the forward velocity remain similar among the top three performing profiles at a given wavelength, a marginal increase in velocity is observed with an increasing wavelength. Importantly, the optimal profiles outperform both conventional profiles in terms of forward velocity across the entire range of frequencies and wavelengths considered in Fig. 10. In addition to higher propulsion speed, the optimal profiles also demonstrate reduced work consumption compared to the anguilliform and carangiform profiles, thereby offering a dual advantage in achieving higher propulsive efficiency.
A similar increasing trend is observed in the work consumption of the undulating foil, where the required energy rises with frequency and shows a modest increase with wavelength. The efficiency trends indicate that the best-performing optimal profiles demonstrate superior performance across all wavelengths compared to the conventional profiles, with efficiency consistently increasing as the frequency rises, highlighting the pivotal role of frequency in enhancing propulsive efficiency. A particularly notable jump in efficiency is observed at $\lambda^{*}=0.50$ for $f=0.5$, whereas at $\lambda^{*}=0.6$ and $0.7$, the efficiency shows a smoother and more gradual increase with frequency. Moreover, the optimized profiles differ significantly from conventional swimming modes, which are limited to a maximum efficiency of approximately 22% and perform optimally only at lower frequencies. In contrast, the optimal profiles not only achieve substantially higher efficiencies, exceeding 80% in some cases, but also maintain high performance over a broader frequency range. These findings highlight the robustness and effectiveness of the optimized designs, demonstrating their superiority in propulsive performance and work consumption efficiency compared to traditional undulatory motions.
Figure 10: Comparison of the mean forward velocity (left column), work consumption (middle column), and propulsive efficiency (right column) as functions of undulating frequency $f$, for the top three optimized swimming profiles at four fixed wavelengths: $\lambda^{*}=0.50$ (top row), 0.60 (second row), and 0.70 (bottom row). The optimal profile (black solid line) consistently outperforms the second-best (blue dashed line) and third-best (red dash-dotted line) profiles in terms of efficiency across all wavelengths and frequencies. Notably, the efficiency trends (right column) highlight the strong frequency dependence, with substantial gains observed at higher frequencies, especially at $\lambda^{*}=0.50$. The results also show increasing work requirements at higher frequencies (middle column), while maintaining or improving forward velocity (left column), demonstrating the balance achieved by the optimal profiles between thrust and energy expenditure.
4.2 Wake Topology
@@
The wake topology of the optimal profile, along with those of the anguilliform and carangiform profiles for comparison at $f=4$ and $\lambda^{*}=0.5$, is presented in Fig. 11 for over a complete oscillating cycle. Although the global wake structures appear qualitatively similar across all three cases, a distinct difference is evident in the vortex roll-up behavior. The optimal profile exhibits a more pronounced and coherent rolling of vortices shed from both the lower and upper surfaces, particularly in the mid-to-posterior region of the body. This feature is relatively weaker in the anguilliform profile and virtually absent in the carangiform case. This enhanced vortex roll-up in the optimal configuration can be linked to the unique kinematic features identified in the backbone motion (see Fig. 9). The optimal profile exhibits reduced tail excursion and an out-of-phase head motion relative to the anguilliform profile. These features modify the pressure distribution and vorticity generation along the body, promoting stronger and more organized vortex formation. Consequently, the resulting wake pattern in the optimal case supports more efficient momentum transfer to the fluid, which aligns with the significant reduction in the work required to undulate the body and the corresponding increase in propulsive efficiency observed in the performance metrics.
@@
Figure 12 presents the instantaneous pressure contours for one undulatory cycle for the optimal profile, shown alongside the anguilliform and carangiform profiles for a comparative analysis. Distinct differences in the pressure distribution are evident among the three profiles, particularly near the anterior and posterior regions. These variations arise due to the out-of-phase head motion exhibited by the optimal profile compared to the anguilliform case. The overall pressure magnitude is higher in the optimal profile, which indicates stronger flow-body interaction and better hydrodynamics performance. In contrast, the anguilliform profile displays relatively lower pressure levels, while the carangiform profile exhibits minimal pressure variation near the head. The stronger flow–body interaction in the optimal profile is further supported by the hydrodynamic performance metrics shown in Fig. 13, which presents the instantaneous variations in axial force, lateral force, and the work performed by the foil over one undulatory cycle. The optimal profile exhibits significantly higher force magnitudes with minimal phase lag. A particularly notable observation is that the instantaneous work performed by the optimal profile becomes negative during certain phases of the cycle, a feature not observed in the carangiform profiles. This negative work indicates that the surrounding fluid imparts energy to the body, effectively performing work on the foil. Such fluid-to-body energy transfer facilitates partial energy recovery and reduces the net energetic cost of undulation, thereby enhancing overall propulsive efficiency.
Additionally, a 3D simulation of the optimal profile is performed at $f=4.5$ and $\lambda^{*}=0.5$ to verify the persistence of its superior propulsive efficiency. This case is intended to assess whether the observed performance is intrinsic to the profile or merely a consequence of 2D effects. The 3D model is constructed as an extension of the 2D configuration in the spanwise ($z$) direction, with a spanwise length of $6L$ and periodic boundary conditions imposed along that direction, effectively modeling an infinitely long foil massey2024resolvent . Although this setup is not representative of real biological swimmers such as eels, jackfish, or harbor seals fardi2025characterizing , it serves as a validation benchmark for the 2D optimal case. The computational grid consists of $512\times 400\times 120$ points in the radial, angular, and spanwise directions, respectively. A comparison of hydrodynamic metrics between the 2D and 3D configurations is shown in Fig. 14, while iso-surfaces of $z$-vorticity at five time instants are presented in Fig. 15. The results demonstrate close agreement, confirming the persistence of efficiency in three dimensions and further establishing the validity of the two-dimensional simulations.
@@
Figure 11: Instantaneous vorticity contours, ranging from -10 to 10, illustrating the wake topology for the optimal ($\eta$=65%), anguilliform ($\eta$=14%), and carangiform ($\eta$=13%) swimming profiles.
@@
Figure 13: Instantaneous variations of hydrodynamic quantities over three undulatory cycles for the optimal, anguilliform, and carangiform profiles: (a) axial force, (b) lateral force, and (c) instantaneous work performed by the foil. The optimal profile (solid red line) exhibits significantly higher axial and lateral force magnitudes with minimal phase lag. The instantaneous work in the optimal case becomes negative during one-half portions of the cycle, suggesting energy recovery from the fluid. Such an effect is absent in the anguilliform (green dash-dot line) and carangiform (blue dash-dot-dot line) profiles.
Figure 14: Comparison of hydrodynamic metrics, (a) lift and drag coefficients and (b) work consumption, between two- and three-dimensional simulations of the optimal profile. The results confirm the persistence of propulsive efficiency in 3D and additionally serve as validation of the two-dimensional simulations.
Figure 15: Iso-surfaces of spanwise vorticity ($z$-vorticity) at five time instants for the optimal profile at $f=4.5$ and $\lambda^{*}=0.5$. The left column shows angled 3D views of the vortex structures, while the right column presents their corresponding projections in the xyxy-plane.
@@ 4.3 Force Characteristics
It becomes essential to investigate the distribution of flow-induced forces in greater detail, particularly how these forces contribute to propelling the body. The interaction between these surface forces and the local flow velocity governs the instantaneous energy exchange and determines the overall energy expenditure. Consequently, a detailed analysis of the spatial and temporal distribution of work along the undulating body is critical for elucidating the mechanisms that lead to high propulsive efficiency. To this end, the axial force comprising both pressure and shear stress contributions is first decomposed into local thrust and drag components, corresponding to the positive and negative values of the local axial force, respectively. These components are further resolved into push and pull forces du2019thrust , based on positive and negative pressure relative to the ambient, and are visualized using directional vectors, as illustrated in Fig. 16, at five representative time instants over a complete undulatory cycle. From this point onward, the comparison is restricted to the optimal and anguilliform profiles, for $f=4.0$ and $\lambda^{*}=0.5$, as the optimal profile shares greater kinematic similarity with the anguilliform mode.
@@
In the optimal profile, the anterior region along with the subsequent 15–30% of the body length experiences a relatively stronger thrust pull compared to the anguilliform mode at several time instants during the undulatory cycle. In addition to the observed thrust pull, the anterior region of the optimal profile also encounters additional drag forces during certain phases of the undulatory cycle. Since the motion from the midbody to the posterior region is largely similar between the optimal and anguilliform profiles, no significant qualitative differences are observed in the spatial distribution of thrust and drag components or the corresponding push and pull forces. However, the magnitudes of all four force components are noticeably higher in the optimal profile, specifically in the 66-95% region, and thus, lead to faster propulsion. This comparison is also illustrated in Fig. 18a, where the bar graphs demonstrate the enhanced force amplitudes associated with the optimized motion. Furthermore, the thrust and drag forces in the optimal profile are predominantly governed by push-type contributions, whereas in the anguilliform case, they are primarily influenced by pull-type forces. In other words, the pushing effect is more pronounced in the optimal profile. This distinction is illustrated in Fig. 18b, where the relative proportions of the push and pull forces are presented as bar plots, highlighting the marginally dominant push contribution in the optimized configuration. As a consequence, it is revealed that the anterior and posterior regions of the optimal profile play a pivotal role in its superior performance, exhibiting motion characteristics and force distributions that significantly deviate from conventional swimming modes.
@@
Despite achieving faster propulsion, the optimal profile exhibits reduced work (or energy) consumption associated with body undulation. To investigate this characteristic, a similar decomposition of surface work is performed, distinguishing between positive work (or work-in) and negative work (or recovery). This decomposition, visualized at five representative time instants during an undulating cycle using vertical arrows in Fig. 17, illustrates the distribution of surface work along the body. Outward-pointing arrows indicate work-in, where the foil performs work on the surrounding fluid, while inward-pointing arrows represent recovery, signifying regions where the fluid performs work on the foil, thereby assisting the undulatory motion in terms of energy recovery. Again, it is evident that the anterior and posterior regions are primarily responsible for the differences in work-in and recovery between the profiles. Both the work-in and recovery components are found to be significantly higher in the optimal profile compared to the anguilliform profile. Figure 19a illustrates the integrated magnitudes of these components, computed over 100 time instants within one undulatory cycle. These bar plots clearly show that the optimal profile not only performs greater positive work but also benefits from enhanced energy recovery. Moreover, the relative proportion between work-in and recovery, depicted in Fig. 19b, reveals that the anguilliform profile exhibits a higher ratio of work-in to recovery compared to the optimal profile. It suggests that the optimal profile requires relatively less mechanical effort to sustain its undulatory motion, which in turn contributes to its enhanced propulsive efficiency.
@@
To identify the contribution of different body segments to the work-in and energy recovery during an undulatory cycle, the body is divided into three distinct regions: anterior (0–30% of body length), midbody (30–70%), and posterior (70–100%). The work-in and recovery contributions from each region are then analyzed over a complete undulation cycle. The relative proportions contributing to work-in and energy recovery across the three defined body regions for both the optimal and anguilliform profiles are presented in Fig. 19c using bar graphs. A detailed comparison of the regional energy dynamics reveals distinct performance characteristics across the three body segments. In the anterior region, the optimal profile performs less mechanical work on the surrounding fluid compared to the anguilliform profile, and although the corresponding energy recovery is also lower, the net energy balance is more favorable. It indicates that the optimal profile requires reduced mechanical effort to undulate its anterior portion. In the midbody region, both profiles exhibit similar levels of work-in. However, the anguilliform profile exhibits slightly higher energy recovery, highlighting more efficient energy exchange in this segment. In contrast, the posterior region of the optimal profile demonstrates a substantial increase in energy recovery relative to the anguilliform profile, despite only a marginal increase in work-in. This highlights the posterior segment as a key contributor to energy reutilization and overall propulsive efficiency in the optimal profile. Thus, the optimal profile emerges as an efficient swimmer by leveraging reduced mechanical effort in the anterior region and enhanced energy recovery in the posterior region, collectively contributing to its superior propulsive performance.
@@
Figure 19: (a) Integrated magnitudes of positive (work-in) and negative (recovery) work components within a complete undulatory cycle for the optimal and anguilliform profiles. (b) Corresponding relative percentages of work-in and recovery. (c) Relative proportion of work-in and recovery at three distinct segments of the body: anterior (0–30% of body length), midbody (30–70%), and posterior (70–100%).

### v2

@@ 4.1 Performance Analysis of Optimal Profiles
To identify the optimal swimming profile, we employ the MixMOBO algorithm within a six-dimensional design space. The objective of the optimization is to maximize the propulsive efficiency $\eta$.
The optimization process begins with 50 randomly generated initial samples, followed by 150 optimization iterations (epochs) guided by the MixMOBO strategy. Figure 9 illustrates the evolution of the best-obtained propulsive efficiency over the optimization cycle, together with the corresponding optimal swimming profiles and undulation kinematics. Key parameters of several high-performing cases are summarized in Table 5. The optimized profiles achieve efficiencies in the range of $52\%$$54\%$, indicating a noticeable enhancement of roughly $24\%$$28\%$ compared to the maximum efficiency ($\sim 42\%$) attained by the reference carangiform and anguilliform swimming modes within the same kinematic parameter space. Furthermore, the best-performing optimized profiles attain their maximum propulsive efficiency at low undulation frequencies, consistent with the trends observed for conventional carangiform and anguilliform swimming modes. Notably, the optimal wavelength values of these profiles fall within the high-efficiency wavelength range characteristic of anguilliform swimming, as shown in Fig. 8.
It is important to emphasize that the ranking of the optimized profiles reported in Table 5 is determined by their peak propulsive efficiency at specific wavelength–frequency combinations. Accordingly, the profiles are locally optimal under their respective kinematic conditions. However, such a ranking does not fully characterize the overall performance of these profiles across the entire frequency–wavelength domain. To obtain deeper insight into their global performance characteristics, additional simulations are performed for the four best optimized profiles across the frequency–wavelength parameter space. The resulting propulsive efficiency contours, presented in Fig. 10, provide a comprehensive assessment of the performance of each optimized profile beyond its locally optimal operating point.
It is observed that the optimized profile (denoted as the $1^{\mathrm{st}}$ best profile in Table 5), hereafter referred to as the optimal profile, attains a maximum propulsive efficiency of approximately $57\%$ at a low undulation frequency of $f=0.2$. Moreover, the optimal profile sustains a high efficiency level, ranging from $49\%$ to $57\%$ (corresponding to an enhancement of roughly $16\%$$35\%$ relative to the baseline swimming modes), over a broad region of the frequency–wavelength parameter space, indicating robust performance across a wide range of kinematic conditions. All other optimized profiles exhibit qualitatively similar efficiency trends, with elevated performance concentrated at low frequencies and moderate wavelengths; however, their peak efficiencies are comparatively lower and the regions of sustained high performance are more limited. Based on this global assessment across the entire frequency–wavelength parameter space, the optimized profiles are hereafter ranked accordingly.
It should also be noted that the present study considers an idealized representation of body kinematics. In real fish, physiological and biomechanical constraints may limit the ability to achieve the optimized body profiles identified in this work. Furthermore, the present analysis is based on two-dimensional simulations; therefore, three-dimensional swimming effects, such as spanwise flow structures and body–fin interactions, are not accounted for and may influence the optimal swimming characteristics.
@@
To investigate the differences in the kinematic behavior between the optimized and conventional swimming modes, the backbone (centerline) undulation of the NACA 0012 airfoil over six time instances within one oscillation cycle is presented in Fig. 11. The optimized profiles exhibit similar undulatory patterns, with the main distinction being a more pronounced head motion in the optimal profile. Compared to the anguilliform profile, the optimal profile aligns closely over a significant region that spans approximately 30% to 87% of the body length. However, at the leading and trailing edges, notable differences are evident. The optimal profile exhibits head motion in the opposite direction to that of the anguilliform profile, and the tail amplitude is marginally lower, with the anguilliform tail reaching a maximum lateral amplitude near 0.1. These distinctive features, specifically the out-of-phase head motion, may contribute to the superior propulsive performance exhibited by the optimized profiles. The observed increase in efficiency is likely associated with these deviations from the anguilliform pattern in both head and tail kinematics. To further examine this hypothesis, a detailed analysis of the wake structures and force distributions is presented in the following discussion.
@@
Figure 9: Convergence history of the MixMOBO algorithm illustrating the evolution of the best-obtained propulsive efficiency over 200 search iterations. The inset shows the corresponding optimal swimming profile shapes along with the undulation kinematics of the NACA 0012 airfoil.
Name Design Epoch $\omega_{1}$ $\omega_{2}$ $\omega_{3}$ $\omega_{4}$ $\omega_{5}$ $\lambda^{*}$ $f$ $\eta$%
$opt_{1}$ 4th Best 50 -0.866 0.250 0.187 0.041 0.108 1.1 0.2 44.75
$opt_{2}$ 3rd Best 57 -0.866 0.250 0.187 -0.041 0.108 0.9 1.0 47.45
$opt_{3}$ 2nd Best 83 -0.866 0.250 0.187 -0.062 0.433 0.9 0.2 47.77
$opt_{4}$ 1st Best 105 -0.866 0.0 0.0 0.0 0.500 1.3 0.2 51.08
@@
Design Epoch $\omega_{1}$ $\omega_{2}$ $\omega_{3}$ $\omega_{4}$ $\omega_{5}$ $\lambda^{*}$ $f$ $\eta$%
4th Best 143 -0.8660 0.2500 0.1875 -0.0406 0.1083 1.3 0.2 52.87
3rd Best 130 -0.8660 0.0 0.2165 -0.0312 0.2165 1.2 0.2 53.57
2nd Best 141 -0.8660 0.0 0.2165 0.0 0.2500 1.3 0.2 53.81
1st Best 160 -0.866 0.2500 0.1875 0.0406 0.1083 1.3 0.2 54.76
Figure 10: Contour map of the propulsive efficiency in the frequency–wavelength space for (a) $1^{st}$ best, (b) $2^{nd}$ best, (c) $3^{rd}$ best, and (d) $4^{th}$ best profiles at $\mbox{Re}=1000$.
@@
Figure 11: Time evolution of the backbone (centerline) kinematics of the top three optimized swimming profiles compared with the conventional anguilliform and carangiform modes at five different phases within one undulation cycle ($t/T=0.0$, $0.2$, $0.4$, $0.6$, $0.8$, and $1.0$). The optimized profiles exhibit notable differences in head and tail motions compared to the anguilliform profile.
@@
The wake topology of the optimal profile, along with those of the anguilliform and carangiform profiles for comparison at $f=0.2$ and $\lambda^{*}=1.1$, is presented in Fig. 12 for over a complete oscillating cycle. Although the global wake structures appear qualitatively similar across all three cases, a distinct difference is evident in the vortex roll-up behavior. The optimal profile exhibits a more pronounced and coherent rolling of vortices shed from both the lower and upper surfaces, particularly in the mid-to-posterior region of the body. This feature is relatively weaker in the anguilliform profile and virtually absent in the carangiform case. This enhanced vortex roll-up in the optimal configuration can be linked to the unique kinematic features identified in the backbone motion (see Fig. 11). The optimal profile exhibits reduced tail excursion and an out-of-phase head motion relative to the anguilliform profile. These features modify the pressure distribution and vorticity generation along the body, promoting stronger and more organized vortex formation. Consequently, the resulting wake pattern in the optimal case supports more efficient momentum transfer to the fluid, which aligns with the significant reduction in the work required to undulate the body and the corresponding increase in propulsive efficiency observed in the performance metrics.
@@
Figure 13 presents the instantaneous pressure contours for one undulatory cycle for the optimal profile, shown alongside the anguilliform and carangiform profiles for a comparative analysis. Distinct differences in the pressure distribution are evident among the three profiles, particularly near the anterior and posterior regions. These variations arise due to the out-of-phase head motion exhibited by the optimal profile compared to the anguilliform case. The overall pressure magnitude is higher in the optimal profile, which indicates stronger flow-body interaction and better hydrodynamics performance. In contrast, the anguilliform profile displays relatively lower pressure levels, while the carangiform profile exhibits minimal pressure variation near the head. The stronger flow–body interaction in the optimal profile is further supported by the hydrodynamic performance metrics shown in Fig. 14, which presents the instantaneous variations in axial force, lateral force, and the work performed by the foil over one undulatory cycle. The optimal profile exhibits significantly higher force magnitudes with minimal phase lag. Although the mean work consumption of the optimal profile is slightly higher than that of the conventional profiles, it achieves faster propulsion, resulting in improved overall efficiency. In other words, the optimizer identifies a profile that prioritizes higher propulsive speed. The contour maps of mean work and propulsive speed for both the optimal and anguilliform profiles are presented in Fig. 15. It is observed that the optimal profile achieves a higher propulsive speed than the anguilliform case, albeit at the expense of increased work consumption, as illustrated in the region $0.9\leq\lambda^{*}\leq 1.5$ in Fig. 15, where enhanced efficiency is observed. Importantly, the relative increase in propulsive speed slightly exceeds the corresponding increase in work, indicating a more favorable performance trade-off. Consequently, this imbalance leads to an overall improvement in propulsive efficiency, rendering the optimal profile superior to the anguilliform profile.
@@
Figure 12: Instantaneous vorticity contours, ranging from -1 to 1, illustrating the wake topology for the optimal ($\eta$=57%), anguilliform ($\eta$=42%), and carangiform ($\eta$=36%) swimming profiles.
@@
Figure 14: Instantaneous variations of hydrodynamic quantities over three undulatory cycles for the optimal, anguilliform, and carangiform profiles: (a) axial force, (b) lateral force, and (c) instantaneous work performed by the foil. The optimal profile (solid red line) exhibits significantly higher axial and lateral force magnitudes with minimal phase lag.
Figure 15: Comparison of contour maps for the time-averaged propulsive velocity (top row) and mean work consumption (bottom row) across the frequency–wavelength domain. Results for the optimal swimming profile are shown in the left column, and those for the conventional anguilliform profile are shown in the right column.
@@ 4.2 Force Characteristics
It becomes essential to investigate the distribution of flow-induced forces in greater detail, particularly how these forces contribute to propelling the body. The interaction between these surface forces and the local flow velocity governs the instantaneous energy exchange and determines the overall energy expenditure. Consequently, a detailed analysis of the spatial and temporal distribution of work along the undulating body is critical for elucidating the mechanisms that lead to high propulsive efficiency. To this end, the axial force comprising both pressure and shear stress contributions is first decomposed into local thrust and drag components, corresponding to the positive and negative values of the local axial force, respectively. These components are further resolved into push and pull forces (Du Clos et al., 2019), based on positive and negative pressure relative to the ambient, and are visualized using directional vectors, as illustrated in Fig. 16, at five representative time instants over a complete undulatory cycle. From this point onward, the comparison is restricted to the optimal and anguilliform profiles, for $f=0.2$ and $\lambda^{*}=1.1$, as the optimal profile shares greater kinematic similarity with the anguilliform mode.
@@
In the optimal profile, the anterior region along with the subsequent 15–30% of the body length experiences a relatively stronger thrust pull compared to the anguilliform mode at several time instants during the undulatory cycle. In addition to the observed thrust pull, the anterior region of the optimal profile also encounters additional drag forces during certain phases of the undulatory cycle. Since the motion from the midbody to the posterior region is largely similar between the optimal and anguilliform profiles, no significant qualitative differences are observed in the spatial distributions of thrust and drag components or in the corresponding push and pull forces. However, the magnitudes of all four force components are noticeably higher in the optimal profile, specifically in the 66-95% region, and thus, lead to faster propulsion. This comparison is also illustrated in Fig. 18a, where the bar graphs demonstrate the enhanced force amplitudes associated with the optimized motion. Furthermore, the thrust and drag forces in the optimal and anguilliform profiles are predominantly governed by push-and pull-type contributions, respectively. This is illustrated in Fig. 18b, where the relative proportions of the push and pull forces are presented as bar plots, highlighting the marginally dominant push-type thrust contribution in the optimized configuration. As a consequence, it is revealed that the anterior and posterior regions of the optimal profile play a pivotal role in its superior performance, exhibiting motion characteristics and force distributions.
@@
Despite achieving faster propulsion, the optimal profile exhibits reduced relative work (or energy) consumption associated with body undulation. To investigate this characteristic, a similar decomposition of surface work is performed, distinguishing between positive work (or work-in) and negative work (or recovery). This decomposition, visualized at five representative time instants during an undulating cycle using vertical arrows in Fig. 17, illustrates the distribution of surface work along the body. Outward-pointing arrows indicate work-in, where the foil performs work on the surrounding fluid, while inward-pointing arrows represent recovery, signifying regions where the fluid performs work on the foil, thereby assisting the undulatory motion in terms of energy recovery. Again, it is evident that the anterior and posterior regions are primarily responsible for the differences in work-in and recovery between the profiles. Both the work-in and recovery components are found to be significantly higher in the optimal profile compared to the anguilliform profile. Figure 19a illustrates the integrated magnitudes of these components, computed over 100 time instants within one undulatory cycle. These bar plots clearly show that the optimal profile not only performs greater positive work but also benefits from enhanced energy recovery. Moreover, the relative proportion between work-in and recovery, depicted in Fig. 19b, reveals that the anguilliform profile exhibits a higher ratio of work-in to recovery compared to the optimal profile. It suggests that the optimal profile requires relatively less mechanical effort based on its superior performance to sustain its undulatory motion, which in turn contributes to its enhanced propulsive efficiency.
@@
To identify the contribution of different body segments to the work-in and energy recovery during an undulatory cycle, the body is divided into three distinct regions: anterior (0–30% of body length), midbody (30–70%), and posterior (70–100%). The work-in and recovery contributions from each region are then analyzed over a complete undulation cycle. The relative proportions contributing to work-in and energy recovery across the three defined body regions for both the optimal and anguilliform profiles are presented in Fig. 19c using bar graphs. A detailed comparison of the regional energy dynamics reveals distinct performance characteristics across the three body segments. In the anterior region, the optimal profile performs less mechanical work on the surrounding fluid compared to the anguilliform profile, and although the corresponding energy recovery is also lower, the net energy balance is more favorable. It indicates that the optimal profile requires reduced mechanical effort to undulate its anterior portion. In the midbody region, both profiles exhibit similar levels of energy recovery. However, the anguilliform profile exhibits slightly lower work-in, highlighting more efficient energy exchange in this segment by anguilliform profile. In contrast, the posterior region of the optimal profile demonstrates a substantial increase in energy recovery and a decrease in work-in relative to the anguilliform profile. This highlights the posterior segment as a key contributor to energy reutilization and overall propulsive efficiency in the optimal profile. Thus, the optimal profile emerges as an efficient swimmer by leveraging reduced mechanical effort in the anterior region and enhanced energy recovery in the posterior region, collectively contributing to its superior propulsive performance.
@@
Figure 19: (a) Integrated magnitudes of positive (work-in) and negative (recovery) work components within a complete undulatory cycle for the optimal and anguilliform profiles. (b) Corresponding relative percentages of work-in and recovery. (c) Relative proportion of work-in and recovery at three distinct segments of the body: anterior (0–30% of body length), midbody (30–70%), and posterior (70–100%). It is important to note that thicker bars are used to represent the work-in, while thinner bars indicate the recovered work.